You’ve just opened a margin account with $25,740 at your local brokerage firm. You instruct your broker to purchase 450 shares of Landon Golf stock, which currently sells for $88 per share. Suppose the call money rate is 5.5 percent and your broker charges you a spread of 1.25 percent over this rate. You hold the stock for four months and sell at a price of $95 per share. The company paid a dividend of $0.55 per share the day before you sold your stock.

a. What is your total dollar return from this investment? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

b. What is your effective annual rate of return? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)
Answer #1

A) Initial purchase amount = 450*88 = 39600$

Loan amount = 39600-25740 = 13860$

Interest on above loan = 13860*6.75%*4/12 = 311.85$

Dividend received = 450(0.55) = 247.5$

Proceed from sale of stock = 450*95 = 42750$

Dollar return = 42750+247.5-(39600+311.85)

=3085.65$

Rate of return = 3085.65/25740 = 11.9878% per four month

B) Effective annual rate of return = 11.9878*12/4 = 35.9633%